Circular No. 6902

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bu IN NGC 3368 U. Munari, R. Barbon, L. Tomasella, and M. Rejkuba, Asiago Astro-

physical Observatory, communicate: “Echelle CCD spectra of SN 1998bu obtained with the Asiago 1.82-m telescope during May 11.81–11.90 UT reveal strong interstellar Na I D absorption. The Na I absorption from M96 is seen at heliocentric radial velocity +744.8 ± 0.3 km/s, with an equivalent width for the Na I D_{1} of 0.035 nm, which corresponds to E(B − V) = 0.15 according to the Munari and Zwitter (1997, A.Ap. 318, 269) calibration. The Na I absorption arising inside our Galaxy is seen at heliocentric radial velocity −6 ± 1.5 km/s, with a Na I D_{1} equivalent width of 0.019 nm and corresponding E(B − V) = 0.06.”

R. Evans and W. Wild, University of Chicago, report the following photometric magnitudes obtained with the Yerkes 1-m reflector during May 11.13–11.15 UT: B = 13.1, V = 12.7, R = 12.3, I = 12.2 (comparison star HR 1433).

L. Boschini, Bergamo, Italy, reports the following position for SN 1998bu (cf. IAUC 6899) from an unfiltered CCD image obtained on May 10.84 UT with a 0.25-m telescope: α = 10^{h}46^{m}46^{s}.01, δ = +11^{o}50′07′′.5 (equinox 2000.0), mag 12.6. Identical positions have been measured by D. Hanžl, Brno, Czech Republic, and by G. Masi, Ceccano, Italy; Masi’s unfiltered CCD image, obtained on May 10.821, yields mag 12.6.

Visual magnitude estimates: May 11.83 UT, 11.9 (A. Baransky, Kiev, Ukraine); 11.88, 12.0 (K. Hornoch, Lelekovice, Czech Republic); 12.72, 12.2 (B. Monard, Pretoria, South Africa); 12.8, 12.2 (L. Kiss, Szeged, Hungary); 12.82, 12.0 (Baransky).

SAX J1808.4–3658 = XTE J1808–369 W. Percival, P. Charles, E. Kuulkers, L. Miller, and T. Shahbaz, Uni-

versity of Oxford; and R. McLure and J. Dunlop, University of Edinburgh, report: “We performed infrared observations of the proposed optical counterpart (IAUC 6885) of this object (IAUC 6876, 6877, 6878) with the U.K. Infrared Telescope on May 2.6 and 4.6 UT. We find K = 17.17 and 17.33 (±0.04), respectively, indicating that the source has faded by ∼ 3.5 mag compared to its discovery magnitude. This dramatic fading in the infrared, together with the near-simultaneous abrupt decline in x-rays (IAUC 6897), strengthens its identification as the optical counterpart.”

1998 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green